OpenStudy (anonymous):

State if true or false . A {1,2} B {3,4} A*(B∩∅) = ∅.

OpenStudy (unklerhaukus):

what do you mean by '*' ?

OpenStudy (anonymous):

MULTIPLIED

OpenStudy (anonymous):

@waterineyes you know this ?

OpenStudy (unklerhaukus):

\[A\times(B∩∅) \]\[=A\times∅\]\[=\]

OpenStudy (anonymous):

∅ .. so true right @UnkleRhaukus ?

OpenStudy (unklerhaukus):

im not sure what happens when you multiply a set by the null set,

OpenStudy (anonymous):

that my teacher taught is a ∅ , i wern't sure about what happens with B∩∅ .. SO THANKS UNKLE :)

OpenStudy (unklerhaukus):

actually i think an empty product is equal to one \[A\times\emptyset=1\]

OpenStudy (anonymous):

really ?